OpenStudy (anonymous):

using double-angle or half-angle identities, find the exact value of: sin 5(Pie)/12

OpenStudy (anonymous):

75 degrees?

OpenStudy (anonymous):

No that's not it

OpenStudy (anonymous):

are u sure its limited only for half and double angles?

OpenStudy (anonymous):

yup

OpenStudy (paxpolaris):

\[\sin \left( 5 \pi \over 12 \right) = \sin \left({1 \over 2} \times{5 \pi \over 6} \right)\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

what did you get?

OpenStudy (paxpolaris):

\[\sin \left(5 \pi \over 12 \right) = \[\sqrt{1-\cos \left( 5\pi \over 6 \right)\over2}=\sqrt{1+{\sqrt{3}\over2}\over2}= {{\sqrt{2+\sqrt{3}}\over 2} }=\]

OpenStudy (paxpolaris):

\[={\sqrt{2}+\sqrt{6}\over4}\]